OpenStudy (sleepyjess):

Did I do something wrong on this question? What is the volume of 3.5 moles of oxygen gas at standard temperature and pressure?

OpenStudy (sleepyjess):

\(\sf V=\dfrac{nRT}{P}\\\dfrac{3.5(.0821)(300)}{.97}\)

OpenStudy (sleepyjess):

Came out with 88.8711

OpenStudy (sleepyhead314):

I don't think standard temp is 300

OpenStudy (sleepyjess):

oohh I was looking at how they did the example problem, okie, so it would be \(\sf\dfrac{3.5(.0821)(273)}{1}\)

OpenStudy (sleepyjess):

wheee and that came out to an answer choice :D

OpenStudy (sleepyhead314):

r = 8.31?

OpenStudy (sleepyjess):

8.31???

OpenStudy (sleepyhead314):

instead of 0.0821

OpenStudy (sleepyjess):

why 8.31 though?

OpenStudy (sleepyhead314):

wait

OpenStudy (sleepyhead314):

I was looking at different units xP you're right nvm

OpenStudy (sleepyjess):

okie ^_^

OpenStudy (sleepyjess):

thank you :D